OpenStudy (anonymous):

Solve 2x2 = -10x. Answer x = 2, x = -5 x = 2, x = 5 x = 0, x = 5 x = 0, x = -57

OpenStudy (ash2326):

@mariahmundy94 did you try this? \[2x^2=-10x\]

OpenStudy (anonymous):

thats the problem.. i didnt put a[ ^]

OpenStudy (ash2326):

I'll give you a hint then you have to try. Bring all the terms to one side

OpenStudy (anonymous):

idk ? 2, -5 ?

OpenStudy (ash2326):

No , first bring all the terms to one side and type the equation here

OpenStudy (anonymous):

0, 5

OpenStudy (ash2326):

@mariahmundy94 we have \[2x^2+10x=0\] \[2x \times x+ 2x \times 5=0\] Did you understand till here

OpenStudy (anonymous):

so did i get it right

OpenStudy (ash2326):

Now you take 2x common out of the terms, and tell what you get?

OpenStudy (anonymous):

0 , -5

OpenStudy (ash2326):

Yeah , but how did you get?